What acceleration will result when a 12 N net force applied to a 3kg object? A 6kg object?


1
Expert's answer
2020-11-23T10:30:10-0500

a) By the definition of the Newton's Second Law of Motion, we have:


"F_{net}=ma,""a=\dfrac{F_{net}}{m}=\dfrac{12\ N}{3\ kg}=4\ \dfrac{m}{s^2}."

b) We can find the acceleration of the 6 kg object using the same formula:


"a=\dfrac{F_{net}}{m}=\dfrac{12\ N}{6\ kg}=2\ \dfrac{m}{s^2}."

Answer:

a) "a=4\ \dfrac{m}{s^2}."

b) "a=2\ \dfrac{m}{s^2}."